OpenStudy (anonymous):

Please explain balancing equation. Especially double bonding. I am confused when to care about the charges and when to put subscriptions and when not to. Please help I only have an hour left.

OpenStudy (paxpolaris):

Balancing equations: there have to be the same number of atoms of each element on both sides. its easier to help on a specific question/example ... I have to go ... but if you have question ... & no one here to answer them post a link in math chat-room saying it's a chem question...

OpenStudy (anonymous):

I understand how to balance an equation but I'm confused because they add coefficients but not subscripts. (I understand why they do that) But my question is, sometimes when balancing equations, a compound's charges does not match. for example in AB+ CD, A has a charge of -1 and B has a charge of +2. Then do we add a subscript of 2 in A or coefficient? because if I add coefficient, then B is also affected!

OpenStudy (anonymous):

I have some discussion about balancing equations here: http://www.chemteam.info/Equations/Equat … Here: http://www.chemteam.info/Equations/WS-Ba

OpenStudy (anonymous):

o balance an equation, there must be equal amounts of atoms on each side of the equation. Look at the UNBALANCED equation of the formation of water: H2 + O2 = H2O The reason this is not balanced is because one oxygen atom is not accounted for on the right side of the equation. You need to add a coefficient somewhere to account for the extra O atom. In the product (H2O) there are twice as many hydrogen atoms as oxygen atoms. So by multiplying the Hydrogen (H2) by 2, we have twice as many hydrogen atoms on both sides of the equation: 2H2 + O2 = H2O. But now there are four hydrogen atoms on one side and two on another, and two oxygen atoms on one side and one on the other. Solution: Multiply H2O by 2. 2H2 + O2 = 2H2O Now the equation is balanced. Hope this helps

OpenStudy (anonymous):

Thanks :).

OpenStudy (paxpolaris):

In the link: http://www.chemteam.info/Equations/Equations.html you can look at Balancing Equations Videos

OpenStudy (anonymous):

Yeah i had that site to but it wouldn't paste :).